OpenStudy (anonymous):

what is 2x + y = 14 in slope intercept form

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

how...

OpenStudy (anonymous):

subtract 2x from both sides

OpenStudy (anonymous):

2x -14? and 2x-2x?

OpenStudy (natalieee):

-2 in the 14 is the intercept

OpenStudy (anonymous):

2x + y = 14 -2x -2x y = -2x + 14

OpenStudy (anonymous):

it that it?

OpenStudy (natalieee):

-2 is the slope and 14 is the intercept

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wouldn't i put -2/1 ?

OpenStudy (natalieee):

yes

OpenStudy (anonymous):

okay, thanks :D

OpenStudy (natalieee):

your welcome (: